# Efficient Batch and Recursive Least Squares for Matrix Parameter Estimation with Application to Adaptive MPC

### v1

@@ Abstract
Traditionally, batch least squares (BLS) and recursive least squares (RLS) are used for identification of a vector of parameters that form a linear model. In some situations, however, it is of interest to identify parameters in a matrix structure. In this case, a common approach is to transform the problem into standard vector form using the vectorization (vec) operator and the Kronecker product, known as vec-permutation. However, the use of the Kronecker product introduces extraneous zero terms in the regressor, resulting in unnecessary additional computational and space requirements. This work derives matrix BLS and RLS formulations which, under mild assumptions, minimize the same cost as the vec-permutation approach. This new approach requires less computational complexity and space complexity than vec-permutation in both BLS and RLS identification. It is also shown that persistent excitation guarantees convergence to the true matrix parameters. This method can used to improve computation time in the online identification of multiple-input, multiple-output systems for indirect adaptive model predictive control.
I Introduction
@@
A significant drawback, however, is that the vec-permutation method increases the dimension of the linear measurement process by using the Kronecker product, introducing extraneous zero terms in the regressor (e.g. equation (15) of [4]). This results in increased computational cost and storage requirements. Other methods for identifying parameters in a matrix structure have been proposed including square root filtering [13], multiinnovations [14, 15], and gradient-based methods [16]. However, these methods do not address whether a least squares cost function is globally minimized or the relationship of the method to standard least squares methods.
This work develops a batch and recursive least squares algorithm for identification of matrix parameters which, under mild assumptions, minimizes the same cost function used in the vec-permutation approach. This method provides a $\mathcal{O}(m^{3})$ times improvement in computational cost and a $\mathcal{O}(m^{2})$ times improvement in storage requirements over vec-permutation, where $m\geq 1$ is the number of columns of the identified parameter matrix. We also show how persistent excitation guarantees convergence of the identified matrix parameters to true matrix parameters, which extends established results for identification of vector parameters [17, 18, 19]. Finally, we show how this method can be used to significantly reduce computation time spent on online identification in predictive cost adaptive control (PCAC) [4].
II The Vec-permutation Approach
@@
where $\operatorname{vec}(\cdot)$ is the column stacking operator, positive-definite $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ is the regularization matrix, $\theta_{0}\in{\mathbb{R}}^{n\times m}$ is an initial estimate of $\theta$, and, for all $k\geq 0$, positive-definite $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ is the measurement weighting matrix.
@@
It then follows that, for all $k\geq 0$, the cost function $J_{k}$, given in (II), can be rewritten as
@@
Propositions 1 and 2 respectively give the batch and recursive least squares methods using vec-permutation to minimize cost function $J_{k}$.
@@
Let $N\geq 0$. For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ be positive definite. Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (II), has a unique minimizer whose vectorization is given by
$\operatorname{vec}\left(\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})\right)=\bar{A}_{N}^{-1}\bar{b}_{N},$ (9)
@@
$\bar{A}_{N}$ $\triangleq\bar{R}+\sum_{i=0}^{N}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{\phi}_{i},$ (10)
$\bar{b}_{N}$ $\triangleq\bar{R}\operatorname{vec}(\theta_{0})+\sum_{i=0}^{N}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{y}_{i},$ (11)
and where, for all $0\leq k\leq N$, $\bar{y}_{k}\in{\mathbb{R}}^{mp}$ and $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$ are defined in (4) and (5), respectively.
Proof.
This result follows directly from Lemma A.3. ∎
Proposition 2.
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let $\bar{\Gamma}_{i}\in{\mathbb{R}}^{mp\times mp}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (II), has a unique minimizer, whose vectorization is denoted as
$\bar{\theta}_{k+1}\triangleq\operatorname{vec}\left(\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})\right).$ (12)
Moreover, for all $k\geq 0$, $\bar{\theta}_{k+1}$ is given recursively by
@@
where $\bar{y}_{k}\in{\mathbb{R}}^{mp}$ and $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$ are defined in (4) and (5), respectively, $\bar{P}_{0}\triangleq\bar{R}^{-1}$, and, for all $k\geq 0$, $\bar{P}_{k}\in{\mathbb{R}}^{mn\times mn}$ is positive definite, and hence nonsingular. Moreover, for all $k\geq 0$, $\bar{P}_{k}$ can be expressed recursively as
@@
This result follows directly from Lemma A.3. ∎
An inefficiency with this method is that the Kronecker product in (5) introduces extraneous zero terms in $\bar{\phi}_{k}$ when $m>1$, resulting in a sparse and higher dimension regressor matrix. However, there is no way to simplify the results of Propositions 1 and 2 as the regularization matrix $\bar{R}$ and measurement weighting matrices $\bar{\Gamma}_{k}$ are not necessarily sparse. The computational complexities of BLS and RLS with vec-permutation are shown in Tables I and II respectively.
III Independent Column Weighting
Next, we assume there exist positive-definite $R_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
$\bar{R}=\begin{bmatrix}R_{1}&\cdots&0_{n\times n}\\ \vdots&\ddots&\vdots\\ 0_{n\times n}&\cdots&R_{m}\end{bmatrix}.$ (16)
Furthermore, assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{1,k},\ldots,\Gamma_{m,k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
$\bar{\Gamma}_{k}=\begin{bmatrix}\Gamma_{1,k}&\cdots&0_{p\times p}\\ \vdots&\ddots&\vdots\\ 0_{p\times p}&\cdots&\Gamma_{m,k}\end{bmatrix}.$ (17)
This corresponds to independent weighting of the columns of $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and $\hat{\theta}-\theta_{0}$ in (II). Then, for all $k\geq 0$ and $\hat{\theta}\in{\mathbb{R}}^{n\times m}$, (II) can be rewritten as
@@ III Independent Column Weighting
The following Lemma shows that minimizing the cost function $J_{k}$, given by (18), can be done by separately minimizing $J_{j,k}$ for all $j=1,\ldots,m$.
Lemma 1.
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer given by
Proof.
Note that, for all $j=1,\ldots,m$, since $J_{j,k}$ is a function of only $\hat{\theta}_{j}$, it follows from Lemma A.3 that $J_{j,k}$ has a unique minimizer. Then, since $J_{k}(\hat{\theta})=\sum_{j=0}^{m}J_{j,k}(\hat{\theta}_{j})$, it follows from (22) that the Lemma holds. ∎
Propositions 3 and 4 respectively give the independent column weighting batch and recursive least squares methods to minimize cost function $J_{k}$ given by (18).
@@
Let $N\geq 0$. For all $0\leq k\leq N$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, $J_{N}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer given by
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{N}(\hat{\theta})=\begin{bmatrix}A_{1,N}^{-1}b_{1,N}&\cdots&A_{m,N}^{-1}b_{m,N}\end{bmatrix},$ (23)
where, for all $j=1,\ldots,m$,
$A_{j,N}$ $\triangleq R_{j}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{j,i}\phi_{i},$ (24)
$b_{j,N}$ $\triangleq R_{j}\theta_{j,0}+\sum_{i=0}^{N}\phi_{i}^{\rm T}\Gamma_{j,i}y_{j,i},$ (25)
and where, for all $0\leq k\leq N$, $y_{j,k}\in{\mathbb{R}}^{p}$ is defined in (20) and $\theta_{j,0}\in{\mathbb{R}}^{n}$ is defined (21).
Proof.
It follows from Lemma A.3 that, for all $j=1,\ldots,m$, $\operatorname*{arg\,min}_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{j,N}(\hat{\theta}_{j})=A_{j,N}^{-1}b_{j,N}$. Hence, Lemma 1 implies (23). ∎
Proposition 4.
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (18), has a unique minimizer, denoted as
$\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta})=\begin{bmatrix}\theta_{1,k+1}&\cdots&\theta_{m,k+1}\end{bmatrix},$ (26)
where, for all $j=1,\ldots,m$, $\theta_{j,k+1}\in{\mathbb{R}}^{n}$ is given by
@@
where $y_{j,k}\in{\mathbb{R}}^{p}$ is defined in (20), $\theta_{j,0}\in{\mathbb{R}}^{n}$ is defined (21), $P_{j,0}\triangleq R_{j}^{-1}$, and, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}\in{\mathbb{R}}^{n\times n}$ is positive definite, and hence nonsingular. Moreover, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}$ can be expressed recursively as
@@
For all $j=1,\ldots,m$, let $\theta_{j,k+1}\triangleq\operatorname*{arg\,min}_{\hat{\theta}_{1}\in{\mathbb{R}}^{n}}J_{j,k}$. It then follows from Lemma A.3 that, for all $j=1,\ldots,m$, (27) and (28) hold. Finally, Lemma 1 implies (23). ∎
An advantage of independent column weighting versus vec-permutation is that no Kronecker product is used, implying that no sparse matrices are introduced. The computational complexities of BLS and RLS with independent column weighting are shown in Tables I and II respectively. Independent column weighting results in a $\mathcal{O}(m^{2})$ times improvement in computational complexity over vec-permutation for both BLS and RLS as well as a $\mathcal{O}(m)$ times improvement in space complexity.
IV Independent Identical Column Weighting
Finally, we now assume there exist positive-definite $R\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
$\bar{R}=\begin{bmatrix}R&\cdots&0_{n\times n}\\ \vdots&\ddots&\vdots\\ 0_{n\times n}&\cdots&R\end{bmatrix}=I_{m}\otimes R.$ (30)
Furthermore, assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
$\bar{\Gamma}_{k}=\begin{bmatrix}\Gamma_{k}&\cdots&0_{p\times p}\\ \vdots&\ddots&\vdots\\ 0_{p\times p}&\cdots&\Gamma_{k}\end{bmatrix}=I_{m}\otimes\Gamma_{k}.$ (31)
This corresponds to independent and identical weighting of the columns of $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and $\hat{\theta}-\theta_{0}$ in (II). Then, it follows from Lemma A.2 that, for all $k\geq 0$, (II) can be rewritten as

### v2

@@ Abstract
Traditionally, batch least squares (BLS) and recursive least squares (RLS) are used for identification of a vector of parameters which form a linear model. In some situations however, it is of interest to identify parameters in a matrix structure. In this case, a common approach is to transform the problem into standard vector form using the vectorization (vec) operator and the Kronecker product, known as vec-permutation. However, the use of the Kronecker product introduces extraneous zero terms in the regressor, resulting in unnecessary additional computational and space requirements. This work derives matrix BLS and RLS formulations which, under mild assumptions, minimize the same cost as the vec-permutation approach. This new approach requires less computational complexity and space complexity than vec-permutation in both BLS and RLS identification. It is also shown that persistent excitation guarantees convergence to the true matrix parameters. This method can used to improve computation time in the online identification of multiple-input, multiple-output systems for indirect adaptive model predictive control.
{IEEEkeywords}
Identification, Modeling, Adaptive Systems, MIMO Systems
1 Introduction
@@
A significant drawback, however, is that the vec-permutation method increases the dimension of the linear measurement process by using the Kronecker product, introducing extraneous zero terms in the regressor (e.g. equation (15) of [4]). This results in increased computational cost and storage requirements. Another approach is to apply standard least squares methods to separately identify the columns of the matrix of parameters [2, p. 102] but this method does not address what cost function is being minimized or the relationship to the vec-permutation approach. Other related methods including square root filtering [13], multiinnovations [14], and gradient-based methods [15] also do not address whether a least squares cost function is globally minimized or the relationship to standard least squares methods.
This work derives a batch and recursive least squares algorithm for identification of matrix parameters which, under the assumption of independent residual error and parameter column weighting, minimizes the same cost function used in the vec-permutation approach. This method provides an $\mathcal{O}(m^{3})$ times improvement in computational complexity and an $\mathcal{O}(m^{2})$ times improvement in storage requirements over vec-permutation, where $m\geq 1$ is the number of columns of the identified parameter matrix. We also show how persistent excitation guarantees convergence of the identified matrix parameters to true matrix parameters, which extends established results for identification of vector parameters [16, 17]. However, we show this improvement in computational complexity may come at the cost of performance if the columns of measurement noise are highly correlated. Finally, we show how this method can be used to significantly reduce computation time spent on online identification in predictive cost adaptive control (PCAC) [4].
2 Vec-permutation Least Squares
@@
where $\operatorname{vec}(\cdot)$ is the column stacking operator, positive-definite (and thus, by definition, symmetric) $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ is the regularization matrix, $\theta_{0}\in{\mathbb{R}}^{n\times m}$ is an initial estimate of $\theta$, and, for all $k\geq 0$, positive-definite $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ is the weighting matrix. If no estimate of the parameters $\theta$ is known, common choices in practice for $\theta_{0}$ are $\theta_{0}=0$ or randomly sampled values. Note that while the regularization term $\operatorname{vec}(\hat{\theta}-\theta_{0})^{\rm T}\bar{R}\operatorname{vec}(\hat{\theta}-\theta_{0})$ results in a biased estimate, this term reduces variance and guarantees that the least squares cost (2) has a unique global minimizer which is useful when few data have been collected [17][18, sec. 7.3].
@@
It then follows that, for all $k\geq 0$, the cost function $J_{k}$, given in (2), can be rewritten as
@@
Propositions 1 gives the vec-permutation approach to minimize cost function $J_{k}$.
@@
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let $\bar{\Gamma}_{k}\in{\mathbb{R}}^{mp\times mp}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let $\bar{R}\in{\mathbb{R}}^{mn\times mn}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (2), has a unique minimizer, whose vectorization is denoted as $\bar{\theta}_{k+1}\triangleq\operatorname{vec}(\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta}))$, which is given by
$\bar{\theta}_{k+1}=\bar{A}_{k}^{-1}\bar{b}_{k},$ (9)
@@
$\bar{A}_{k}$ $\triangleq\bar{R}+\sum_{i=0}^{k}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{\phi}_{i},$ (10)
$\bar{b}_{k}$ $\triangleq\bar{R}\operatorname{vec}(\theta_{0})+\sum_{i=0}^{k}\bar{\phi}_{i}^{\rm T}\bar{\Gamma}_{i}\bar{y}_{i},$ (11)
and where, for all $k\geq 0$, $\bar{y}_{k}\in{\mathbb{R}}^{mp}$ and $\bar{\phi}_{k}\in{\mathbb{R}}^{mp\times mn}$ are defined in (4) and (5), respectively. Furthermore, for all $k\geq 0$, $\bar{\theta}_{k+1}$ is given recursively by
@@
where $\bar{P}_{0}\triangleq\bar{R}^{-1}$, and, for all $k\geq 0$, $\bar{P}_{k}\in{\mathbb{R}}^{mn\times mn}$ is positive definite, and hence nonsingular. Moreover, for all $k\geq 0$, $\bar{P}_{k}$ can be expressed recursively as
@@
Since $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (2), is a standard least squares cost with vector parameters, this result follows directly from [3]. ∎
Equations (9) through (13) give the batch least squares solution using vec-permutation while (12) through (14) give the recursive least squares solution using vec-permutation. An inefficiency with this method is that the Kronecker product in (5) introduces extraneous zero terms in $\bar{\phi}_{k}$ when $m>1$, resulting in a sparse and higher dimensional regressor matrix. However, the results of Proposition 1 cannot be simplified since the regularization matrix $\bar{R}$ and weighting matrices $\bar{\Gamma}_{k}$ are not necessarily sparse.
3 Column-by-Column Least Squares
To simplify the vec-permutation approach, we make the assumption that there exist positive-definite $R_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
$\bar{R}=\operatorname{diag}(R_{1},\ldots,R_{m}).$ (15)
Furthermore, we assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{1,k},\ldots,\Gamma_{m,k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
$\bar{\Gamma}_{k}=\operatorname{diag}(\Gamma_{1,k},\ldots,\Gamma_{m,k}).$ (16)
This corresponds to independent weighting of the columns of the residual error, $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and of the regularization term, $\hat{\theta}-\theta_{0}$, in (2). Then, for all $k\geq 0$ and $\hat{\theta}\in{\mathbb{R}}^{n\times m}$, (2) can be rewritten as
@@ 3 Column-by-Column Least Squares
Propositions 2 show that, under assumptions (15) and (16), the cost function $J_{k}$, given by (17), can be minimized by separately updating the columns of the parameter estimate. We call this the column-by-column approach.
@@
For all $k\geq 0$, let $\phi_{k}\in{\mathbb{R}}^{p\times n}$, let $y_{k}\in{\mathbb{R}}^{p\times m}$, and let ${\Gamma}_{1,k},\ldots,{\Gamma}_{m,k}\in{\mathbb{R}}^{p\times p}$ be positive definite. Furthermore, let $\bar{\theta}_{0}\in{\mathbb{R}}^{n\times m}$ and let ${R}_{1},\ldots,R_{m}\in{\mathbb{R}}^{n\times n}$ be positive definite. Then, for all $k\geq 0$, $J_{k}\colon{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$, defined in (17), has a unique minimizer, whose columns are denoted as
$\begin{bmatrix}\theta_{1,k+1}&\cdots&\theta_{m,k+1}\end{bmatrix}\triangleq\operatorname*{arg\,min}_{\hat{\theta}\in{\mathbb{R}}^{n\times m}}J_{k}(\hat{\theta}),$ (22)
which, for all $j=1,\ldots,m$, are given by
$\theta_{j,k+1}=A_{j,k}^{-1}b_{j,k},$ (23)
where
$A_{j,k}$ $\triangleq R_{j}+\sum_{i=0}^{k}\phi_{i}^{\rm T}\Gamma_{j,i}\phi_{i},$ (24)
$b_{j,k}$ $\triangleq R_{j}\theta_{j,0}+\sum_{i=0}^{k}\phi_{i}^{\rm T}\Gamma_{j,i}y_{j,i},$ (25)
and where, for all $k\geq 0$, $y_{j,k}\in{\mathbb{R}}^{p}$ is defined in (19) and $\theta_{j,0}\in{\mathbb{R}}^{n}$ is defined (20). Furthermore, for all $k\geq 0$ and $j=1,\ldots,m$, $\theta_{j,k+1}\in{\mathbb{R}}^{n}$ is given recursively by
@@
where $P_{j,0}\triangleq R_{j}^{-1}$, and, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}\in{\mathbb{R}}^{n\times n}$ is positive definite, and hence nonsingular. Moreover, for all $k\geq 0$ and $j=1,\ldots,m$, $P_{j,k}$ can be expressed recursively as
@@
Note that, for all $j=1,\ldots,m$, $J_{j,k}$ is a function of only $\hat{\theta}_{j}$ and is a standard least squares cost with vector parameters. It then follows from [3] that, for all $j=1,\ldots,m$, $J_{j,k}$ has a unique minimizer, denoted as $\theta_{j,k+1}\triangleq\operatorname*{arg\,min}_{\hat{\theta}_{j}\in{\mathbb{R}}^{n}}J_{j,N}(\hat{\theta}_{j})$, which is given by (23). It also follows from [3] that (24) through (28) hold. Finally, it follows from (17) and (21) that
and (22) follows. ∎
An advantage of the column-by-column approach versus vec-permutation is that no Kronecker product is used, implying that no sparse matrices are introduced. Note that the column-by-column approach can also be derived by applying standard least squares methods to the $m$ columns of (1) [2, p. 102]. However, our derivation further shows the connection to vec-permutation and how the column-by-column approach implicitly implies independent weighting to the columns of the residual error and parameter regularization.
4 Matrix Update Least Squares
Finally, we make the stronger assumption that there exists positive-definite $R\in{\mathbb{R}}^{n\times n}$ such that $\bar{R}$ is block diagonal of the form
$\bar{R}=\operatorname{diag}(R,\ldots,R)=I_{m}\otimes R.$ (29)
Furthermore, we assume that, for all $k\geq 0$, there exist positive-definite $\Gamma_{k}\in{\mathbb{R}}^{p\times p}$ such that $\bar{\Gamma}_{k}$ is block diagonal of the form
$\bar{\Gamma}_{k}=\operatorname{diag}(\Gamma_{k},\ldots,\Gamma_{k})=I_{m}\otimes\Gamma_{k}.$ (30)
This corresponds to independent and identical weighting of the columns of the residual error, $y_{i}-\phi_{i}\hat{\theta}$, $i=0,\ldots,k$, and of the regularization term, $\hat{\theta}-\theta_{0}$, in (2). Then, it follows from Lemma A.1 that, for all $k\geq 0$, (2) can be rewritten as